solve a literal equation for one of the variables:
1. solve for y:2x+3y=9
2. solve for x: 5x+4y+20=0
3. solve the formula for L:P=2L+2W
4. solve the formula for R: E=IR
simply expression with rational exponets
5. simplify: (b with neg.2/3 power)neg.6th exponet
6.simplify: square root of 16a to the 4th power

Answer by stanbon(75887)

1. solve for y:2x+3y=9
y = (-2/3)x+3
---------------
2. solve for x: 5x+4y+20=0
x = (-4/5)y - 4
---------------
3. solve the formula for L:P=2L+2W
L = (P-2W)/2
--------------------
4. solve the formula for R: E=IR
R = E/I
-----------------
simply expression with rational exponets
5. simplify: (b with neg.2/3 power)neg.6th exponet
b^(-2/3)^(-6) = b^(4)
-------------------------
6.simplify: square root of 16a to the 4th power
sqrt(16a^4) = 4a^2
========================
Cheers,
Stan H.